# In the Picture: Medical Imaging Datasets, Artifacts, and their Living Review

### v1

@@ Appendix B Clinical case studies
Chest X-rays are the most commonly performed radiologic examinations worldwide (UNSCEAR et al., 2000), requiring significant expertise for accurate and meaningful interpretation (Van Ginneken et al., 2001). The advent of artificial intelligence is a game changer for automatic chest X-ray diagnosis, marked by the release of the large-scale NIH-CXR14 dataset (Wang et al., 2017). This milestone spurred a wave of studies, including claims of machine learning models achieving radiologist-level performance, exemplified by CheXNet’s pneumonia detection model (Rajpurkar et al., 2017). However, these claims that have been criticized for relying on shortcuts, such as the use of chest drains for pneumothorax classification (Oakden-Rayner et al., 2020; Jiménez-Sánchez et al., 2023), and for demonstrating low inter-rater agreement for pathologies like pneumonia (Damgaard et al., 2023). The latter may be attributed to the fact that pneumonia is a differential diagnosis requiring clinical information – laboratory tests, physical examinations, symptoms and signs, etc. – that extends beyond chest X-ray images (Bustos et al., 2020). Subsequently, additional datasets such as CheXpert (Irvin et al., 2019), MIMIC-CXR (Johnson et al., 2019), and PadChest (Bustos et al., 2020) have become widely used in the research community.
@@ Appendix C Living review
In Section 2, we propose the living review to keep track of medical imaging datasets and their related research artifacts. The resources that we consider as research artifacts are broadly classified into two categories, see Table C1:
@@
Our living review consists of three parts, as illustrated in Figure 2:
@@
Our demo is vailable at http://130.226.140.142. We show a screenshot of the demo in Figure C1.
@@
Research artifact Citation f. Resource type Example
@@
Figure C2. A screenshot of our SQL database in DBeaver.

### v2

@@ Appendix B Clinical case studies
We present four case studies, illustrated in Fig. 3, with key points highlighted in Table C4.
@@
Chest X-rays are the most commonly performed radiologic examinations worldwide (UNSCEAR et al., 2000), requiring significant expertise for accurate and meaningful interpretation (Van Ginneken et al., 2001). The advent of artificial intelligence is a game changer for automatic chest X-ray diagnosis, marked by the release of the large-scale NIH-CXR14 dataset (Wang et al., 2017). This milestone spurred a wave of studies, including claims of machine learning models achieving radiologist-level performance, exemplified by CheXNet’s pneumonia detection model (Rajpurkar, 2017). However, these claims that have been criticized for relying on shortcuts, such as the use of chest drains for pneumothorax classification (Oakden-Rayner et al., 2020; Jiménez-Sánchez et al., 2023), and for demonstrating low inter-rater agreement for pathologies like pneumonia (Damgaard et al., 2023). The latter may be attributed to the fact that pneumonia is a differential diagnosis requiring clinical information – laboratory tests, physical examinations, symptoms and signs, etc. – that extends beyond chest X-ray images (Bustos et al., 2020). Subsequently, additional datasets such as CheXpert (Irvin et al., 2019), MIMIC-CXR (Johnson et al., 2019), and PadChest (Bustos et al., 2020) have become widely used in the research community.
@@ Appendix C Living review
In Section 2, we propose a proof of concept for the faliving review to keep track of medical imaging datasets and their related research artifacts. The resources that we consider as research artifacts are broadly classified into two categories, see Table C1:
@@
Our living review consists of three parts, as illustrated in Fig. 2:
@@
Our demo is available at http://130.226.140.142, and a screenshot is shown in Fig. C1.
Technical details
We collect datasets, papers, and their relations into a SQL database. We visualize the database with DBeaver as shown in Fig. C2. The database comprise of three tables: one for datasets, another for papers, and a third one for datasets usages. The datasets table includes details such as the dataset ID, name and modality of the dataset. The papers table contains information on the ID, BibTeX key name, DOI, and arXiv link. The datasets usages table stores the ID, paper_ID, dataset_ID, annotations, shortcuts, and the taxonomy to which the confounder belongs. Example of annotations include chest drains, segmentation masks, and bounding boxes. Example of shortcuts include dark corners and demographic attributes. Our demo is built with PostgreSQL, Python, Streamlit and the st_link_analysis package.
[…]
Visualization of the living review as a graph node, highlighting the relationships for the NIH-CXR14 dataset.
Figure C2. A screenshot of our SQL database in DBeaver.
[…]
A screenshot of our SQL database in DBeaver displaying the three tables: datasets, papers and dataset usages.
@@
Research artifact Citation function Resource type Example
@@
Case study Tasks and Datasets Annotation practices Demographics Shortcuts Data lifecycle
1: Chest X-rays T: Classification of chest X-ray diseases Experience radiologists vs. automatic NLP extraction Self-reported race/ethinicity Chest drains, hospital scanner, radiographic markers
D: NIH-CXR14, CheXpert, PadChest, MIMIC-CXR Skewed towards elder patients
2: Skin lesions T: Classification, segmentation and lesion localization. Skin type, genetic background Dark corners, hair, patches, rulers, ink marking/staining
D: ISIC, HIBA, PAD-UFES-20, HAM10000, Fitzpatrick17k Recent additions of demographics with HIBA and PAD-UFES-2
3: Fetal ultrasound T: 2D plane localization, fetal biometrics estimation Experience of the sonographer might affect the image quality None to date
D: Fetal Planes DB, HC18, ACOUSLIC-AI, FH-FS-AOP
4:Neuroimaging T: Neurogenerative diseases and stroke detection. Poorly reported: in US studies (2010–2020), 77% report sex, but only 10% report race and 4% ethnicity. Standards (e.g. BIDS, preprocessing, conversion from DICOM to NIfTI format)
D: OASIS, LUMIERE, BraTS, Ocana, TCGA-GBM, ATLAS v2.0, ISLES 2022 Recent inclusion of children and Sub-Saharan African populations in BraTS Data sharing platforms like Datalad, NITRC.org, OpenNeuro
Table C4. Summary highlighting various aspects of the four case studies. T: tasks; D: datasets. Note: this table is not exhaustive but summarizes key points discussed in the manuscript.